1. A block slides down a frictionless plane having an inclination of = 15.0 0 .The block starts from rest at the top, and the length of the incline is 2.00 m.

Find;

(a) the acceleration of the block

(b) its speed when it reaches the bottom of the incline. 



1
Expert's answer
2020-09-28T08:06:46-0400

a) There are two forces that act on the block: the force of gravity (or the weight) "mg" and the normal force "N". The force of gravity can be resolved into two perpendicular components ("F_{\shortparallel} = mgsin\theta" and "F_{\bot} = mgcos\theta").




We can find the acceleration of the block from the Newton's Second Law of Motion:


"\sum F_x = ma_x,""mgsin\theta = ma,""a = gsin\theta = 9.8 \ \dfrac{m}{s^2} \cdot sin15^{\circ} = 2.54 \ \dfrac{m}{s^2}."


b) We can find the speed of the block when it reaches the bottom of the incline from the kinematic equation:


"v_f^2 - v_i^2 = 2as,"

here, "v_i = 0" is the initial speed of the block, "v_f" is the final speed of the block when it reaches the bottom of the incline, "a" is the acceleration of the block, "s = 2.0 \ m" is the distance traveled by the block.

Then, from this formula we can calculate the speed of the block when it reaches the bottom of the incline:


"v_f = \sqrt{2as} = \sqrt{2 \cdot 2.54 \ \dfrac{m}{s^2} \cdot 2.0 \ m} = 3.19 \ \dfrac{m}{s}."

Answer:

a) "a = 2.54 \ \dfrac{m}{s^2}."

b) "v_f = 3.19 \ \dfrac{m}{s}."